OpenStudy (aaronandyson):

An employer find that if he increase wages of each worker by $5 and employs 5 workers less,he increases his weekly wage bill from $3150 to $3250,Taking the original weekly wage of each worker as x; obtain an equation in x and then solve it to find the weekly wage of each worker.

OpenStudy (aaronandyson):

@Directrix @dan815

OpenStudy (baru):

original wage : x original no of workers : N original bill: Nx=3150 bill after modifications (N-5)(N+5)=3250 use substitution to solve these two equations

OpenStudy (baru):

*(N-5)(x+5)=3250

OpenStudy (baru):

which part you having a problem with?

OpenStudy (aaronandyson):

@imqwerty

imqwerty (imqwerty):

where are u stuck?

OpenStudy (aaronandyson):

substitution

imqwerty (imqwerty):

ok

imqwerty (imqwerty):

Nx=3150 can u find out N in terms of x

OpenStudy (anonymous):

̸̶̵̴̛͕͇̾̒̏́͐̌͒̊͆̃̉͗͑̑̿̅̄̚͠͡͝͞γ̴̵̶̸̸̶̵̴̛̛͇͕͕͇̄̅̿̑͑͗̉̃͆̊͒̌͐́̏̒̾̾̒̏́͐̌͒̊͆̃̉͗͑̑̿̅̄̚̚͞͝͡͠͠͡͝͞l̴̵̶̸̸̶̵̴̛̛͇͕͕͇̄̅̿̑͑͗̉̃͆̊͒̌͐́̏̒̾̾̒̏́͐̌͒̊͆̃̉͗͑̑̿̅̄̚̚͞͝͡͠͠͡͝͞l̴̵̶̸̸̶̵̴̛̛͇͕͕͇̄̅̿̑͑͗̉̃͆̊͒̌͐́̏̒̾̾̒̏́͐̌͒̊͆̃̉͗͑̑̿̅̄̚̚͞͝͡͠͠͡͝͞ī̴̵̶̸̸̶̵̴̛̛͇͕͕͇̅̿̑͑͗̉̃͆̊͒̌͐́̏̒̾̾̒̏́͐̌͒̊͆̃̉͗͑̑̿̅̄̚̚͞͝͡͠͠͡͝͞Ɉ̴̵̶̸̸̶̵̴̛̛͇͕͕͇̄̅̿̑͑͗̉̃͆̊͒̌͐́̏̒̾̾̒̏́͐̌͒̊͆̃̉͗͑̑̿̅̄̚̚͞͝͡͠͠͡͝͞n̴̵̶̸̸̶̵̴̛̛͇͕͕͇̄̅̿̑͑͗̉̃͆̊͒̌͐́̏̒̾̾̒̏́͐̌͒̊͆̃̉͗͑̑̿̅̄̚̚͞͝͡͠͠͡͝͞ɒ̴̵̶̸̸̶̵̴̛̛͇͕͕͇̄̅̿̑͑͗̉̃͆̊͒̌͐́̏̒̾̾̒̏́͐̌͒̊͆̃̉͗͑̑̿̅̄̚̚͞͝͡͠͠͡͝͞ʜ̴̵̶̸̛͇͕̄̅̿̑͑͗̉̃͆̊͒̌͐́̏̒̾̚͞͝͡͠Ɔ̴̵̶̸̛͇͕̄̅̿̑͑͗̉̃͆̊͒̌͐́̏̒̾̚͞͝͡͠

OpenStudy (aaronandyson):

Well , I boiled it down to: N^2 - 5N - 3150 = 0

OpenStudy (anonymous):

(͡° ͜ʖ ͡°)

imqwerty (imqwerty):

wait

OpenStudy (aaronandyson):

N^2 - 25N - 3150 = 0

imqwerty (imqwerty):

this is a typo-(N-5)(N+5)=3250 @baru corrected it it is-(N-5)(x+5)=3250

OpenStudy (aaronandyson):

O_O

imqwerty (imqwerty):

what

OpenStudy (aaronandyson):

Well,what next?

imqwerty (imqwerty):

did u understand why is it (N-5)(x+5)=3250

OpenStudy (anonymous):

(͡° ͜ʖ ͡°)

OpenStudy (aaronandyson):

N^2 - 25N - 3150 = 0

imqwerty (imqwerty):

why are u sticking to this? it is wrong actually

imqwerty (imqwerty):

lets start over :) ok?

OpenStudy (aaronandyson):

Okay:)

imqwerty (imqwerty):

lets say that total number of workers=N :) and their original wage be =x

OpenStudy (aaronandyson):

so original its Nx = 3150 , right?

imqwerty (imqwerty):

yes :) can separate N from x anyhow in this equation

OpenStudy (aaronandyson):

x = 3150/N :)

imqwerty (imqwerty):

u separated x well this will work but its better to separate N i'll tell afterwards why it is better to separate N :)

OpenStudy (aaronandyson):

N = 3150/x better?:)

imqwerty (imqwerty):

yes :) now tell what will be the total number of ppl and what will be their new wage :)

OpenStudy (aaronandyson):

(n-5)(x+5) = 3250?

OpenStudy (anonymous):

(͡° ͜ʖ ͡°)

imqwerty (imqwerty):

*N yes thats correct :) now our next step is to fin out x and x is our answer

imqwerty (imqwerty):

to find out x we gotta throw N outta the equation

imqwerty (imqwerty):

so we put N=3150/x in the equation to remove the N ok?

OpenStudy (aaronandyson):

yes

imqwerty (imqwerty):

ok so substitute N=3150/x in the second equation

OpenStudy (aaronandyson):

(3150/x - 5)(x+5) = 3250

imqwerty (imqwerty):

yes now u solve this equation to get x

OpenStudy (aaronandyson):

wait lemme solve :)

imqwerty (imqwerty):

ok (:

OpenStudy (aaronandyson):

(3150 - 5x)(x+5) = 3250x

imqwerty (imqwerty):

yes

OpenStudy (aaronandyson):

x^2 - 25x + 3150 = 0

imqwerty (imqwerty):

that will be -x^2

OpenStudy (aaronandyson):

okay:)

OpenStudy (aaronandyson):

-x^2 - 25x + 3150 = 0 -(x^2 + 25x - 3150) = 0

imqwerty (imqwerty):

go on

OpenStudy (aaronandyson):

I don't know how to factoroise this type of equations:(

imqwerty (imqwerty):

its ok u can use this formula to get x \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a}\] a=coeff of x^2 b=coeff of x c=constant

OpenStudy (aaronandyson):

actually,I can only use the formula when said so in the questions...

OpenStudy (aaronandyson):

eg :"correct the answer to two decimal places" then I HAVE TO use the formula...not otherwise.

OpenStudy (aaronandyson):

its 45?

imqwerty (imqwerty):

yes

imqwerty (imqwerty):

now if we wuld have taken x in terms of N then we wuld have find N and then we wuld have put this value of N in one of the equation to get x so this method involved an extra step this is reason y we didn't took x in terms of N

OpenStudy (aaronandyson):

I know:)

OpenStudy (aaronandyson):

Anyways thanks :) :D B)

imqwerty (imqwerty):

np (B

OpenStudy (baru):

xD...sorry. i should have been more careful :)